### v2

@@ Abstract
This paper introduces ExpertLongBench, an expert-level benchmark containing 11 tasks from 9 domains that reflect realistic expert workflows and applications. Beyond question answering, the application-driven tasks in ExpertLongBench demand long-form outputs that can exceed 5,000 tokens and strict adherence to domain-specific requirements. Notably, each task in ExpertLongBench includes a rubric, designed or validated by domain experts, to specify task requirements and guide output evaluation. Furthermore, we propose Clear, an evaluation framework that supports accurate evaluation of long-form model outputs in our benchmark. To achieve fine-grained, expert-aligned evaluation, Clear derives checklists from both model outputs and references by extracting information corresponding to items in the task-specific rubric. Checklist items for model outputs are then compared with corresponding items for reference outputs to assess their correctness, enabling grounded evaluation. We benchmark 11 large language models (LLMs) and analyze components in Clear, showing that (1) existing LLMs, with the top performer achieving only a 26.8% F1 score, require significant improvement for expert-level tasks; (2) models can generate content corresponding to the required aspects, though often not accurately; and (3) accurate checklist extraction and comparison in Clear can be achieved by open-weight models for more scalable and low-cost usage.
@@ 1 Introduction
There is growing interest in applying large language models (LLMs) to specialized applications that demand domain-specific expertise, such as student tutoring [1], legal case summarization [2], and medical diagnosis [3]. These expert-level tasks pose greater challenges because they require specialized knowledge and strict adherence to domain-specific standards, which are typically acquired through advanced education and professional experience. Furthermore, many real-world expert-level tasks require understanding lengthy inputs [4] and generating complex and nuanced long-form outputs [5, 6].
@@
Existing expert-level benchmarks such as MMLU [7] and GPQA [8] prioritize ease of evaluation of multiple-choice or short-form answers at the expense of alignment with realistic expert applications. Though ExpertQA [9] assesses longer-form, domain-specific outputs, it remains centered on QA scenarios rather than end-to-end expert workflows. In addition to task scope, a key limitation is the lack of evaluation methods tailored to the specific requirements of each task. WildBench [10] creates checklists to evaluate model responses to user queries, yet their model-generated checklists fall short in capturing domain-specific requirements, which often demand expert knowledge to construct. Finally, a broader challenge is the lack of reference outputs in expert-level benchmarks with long-form responses, leading to ungrounded evaluations and oversight of recall estimation.
@@
To address these gaps, this paper first introduces ExpertLongBench, a multi-disciplinary, expert-level benchmark with tasks that require long-form outputs. ExpertLongBench comprises 11 tasks across 9 domains, including law, material science, education, healthcare, chemistry, biology, medicine, finance, and cybersecurity, with 1050 samples in total (Tab. 1). Our benchmark enables more comprehensive assessment of a model’s problem-solving capabilities by including domain-specific end-to-end applications, such as drafting legal briefs or clinical notes, with maximum input and output lengths exceeding 200K and 5K tokens, significantly longer than existing datasets (see comparison in Appendix LABEL:ap:comparison). The completion of these tasks is time-consuming, even for human experts. For example, it takes over 10 hours for a proficient legal practitioner to summarize a complex legal case, as they need to follow the docket and review tens to hundreds of court filings [5]. Moreover, ExpertLongBench provides expert-written reference outputs for each task, facilitating grounded evaluation.
@@
Figure 1: Pipeline of Clear. The example shown is from task T1: multi-document legal case summarization. The checklist mapper takes as input the model output (or human-written reference) and extracts checklist items according to the rubric. Checklists of the model output and the reference are compared at the item level, and the results are subsequently aggregated into the final scores of precision, recall, F1, and accuracy.
@@
To ensure more accurate evaluation of model performance on expert-level tasks, we further propose ClearCheckList-based Expert-level Assessment with Rubric (Fig. 1). We collaborate with domain experts to develop and validate a detailed evaluation rubric for each task. Unlike subjective criteria, such as “usefulness” [9] and “helpful information” [10], our rubric specifies the essential elements required in the output for each application, enabling fine-grained evaluation that closely align with domain-specific requirements. For example, a well-written summary of legal case documents intended for lawyers must accurately identify the cause of action, the relevant statutory or constitutional basis, and the remedy sought. Based on the rubric, Clear breaks down model responses and reference outputs into itemized checklists. For each item, our framework compares the corresponding information extracted from the model’s output against that of the reference output, thereby ensuring a grounded and objective checklist-based assessment. [^1] The benchmark leaderboard and links to code and public data: https://huggingface.co/spaces/launch/ExpertLongBench.
@@
We evaluate the performance of 11 frontier LLMs on ExpertLongBench, including both open-weight and proprietary model families. Our findings are: $\bullet$ The best model achieves an average F1 score of 26.8% across all tasks, revealing that the end-to-end expert-level tasks in ExpertLongBench present significant challenges for existing LLMs. $\bullet$ Despite a low overall output quality, models can generate content that matches over 67% of the aspects required in the checklist, suggesting a risk of models producing content that appears expert-aligned but is potentially misleading due to incorrect generation. $\bullet$ Clear can be configured with open-weight models for cost-effective and scalable benchmarking, as evidenced by a Pearson correlation of 0.88 between the scores assigned by Qwen2.5-72B and GPT-4o for checklist-based evaluation.
@@ 2 Related Work
Previous studies have assessed domain-specific knowledge using multiple-choice questions (e.g., MMLU [7], GPQA [8]) or short-answer questions (e.g., AGIEval [11], SciEval [12], HLE [13], OlympiadBench [14]). Their scope remains constrained to exam-style questions, failing to reflect domain-specific tasks in real world that typically require long-form generation.
ExpertQA curates open-ended questions written by domain experts [9]. However, the answers to these questions are typically around 100 words, as the questions are formulated to address experts’ specific knowledge queries rather than addressing the full scope of an end-to-end workflow. Additionally, the evaluation method relies on human judgments of answer usefulness, which are inherently subjective, coarse-grained, and lack scalability.
ExpertLongBench addresses these gaps by introducing diverse, expert-level tasks that align with complete workflows; incorporating tasks requiring long-form generation and featuring lengthy inputs that are typical in practical use cases; and providing rubric-based evaluation for fine-grained, domain-specific assessment.
Table 1: Benchmark statistics. : rubric is developed by experts; otherwise, it is created by refining and expanding upon established evaluation protocols; : task data is held privately. For each task, we report whether the task data is newly created (✓) or adapted from previous work; the average number of checklist items in each sample (#Rubric); and the average length of the input (#Input) and human reference (#Reference). Several of these tasks feature significantly longer inputs and references compared to existing domain-specific datasets (see Appendix LABEL:ap:comparison).
@@
Beyond traditional shallow n-gram overlap-based metrics [15, 16], BLEURT [17] and BERTScore [18] use contextualized embeddings to assess semantic similarity between generated and reference texts. For more granular evaluation, recent research explores fact decomposition-based methods, extracting atomic facts from generated and reference texts for comparison using NLI models or LLMs. [19, 20]. Nevertheless, the fact decomposition process lacks task-specific considerations for fact granularity, which can result in inconsistent evaluations [21].
Another line of work, commonly termed “LLM-as-a-judge”, prompts LLMs to assess generated outputs against general, high-level criteria such as fluency, coherence, and relevance [22]. To enhance task specificity, checklist-based evaluations are incorporated into WildBench [10] and BiGGenBench [23], which evaluate outputs against instance-specific checklists derived from task requirements. However, the absence of grounding for these checklists in reference outputs can introduce coarse and vague criteria, often leading to subjective assessments and hindering the accurate evaluation of content coverage and relevance.
@@
3 Benchmark Construction with Multi-disciplinary Expert-Level Tasks
3.1 Overview
We introduce ExpertLongBench, a multi-disciplinary benchmark designed to evaluate the capabilities of LLMs on real-world expert-level tasks that require long-form inputs and outputs. ExpertLongBench features a fine-grained evaluation rubric tailored to reflect the rigorous standards and nuanced requirements of expert domains. The benchmark covers 11 expert-level tasks as shown in Tab. 1, including six tasks with newly collected data featuring unique challenges such as super-long input, where experts typically spend over 10 hours to complete [5]. Notably, rubric design is complex and highly time-intensive; for instance, creating the rubric for T1LegalMDS required over 10 hours of expert effort. The selection of these tasks adhere to specific criteria: (1) it is possible to create well-defined rubric for consistent and objective evaluation; (2) it requires domain expertise for both task resolution and assessment; and (3) it is grounded in real-world expert workflows. Addition task selection details are discussed in Appendix A and Appendix B. Generally, each sample in ExpertLongBench includes the following key elements:
  • Human Reference: Each task includes human-authored references in natural language. Details can be referred to Appendix B.
  • Checklist-mapped Reference: A checklist-based, fine-grained evaluation rubric is designed for each task through collaboration with domain experts and by referencing established resources. A corresponding checklist-mapped reference is constructed for every sample, as detailed in §3.2.
Tab. 1 summarizes key statistics of our benchmark. In total, our benchmark contains 1,050 samples, with an average input length and human reference length of 36,204 and 851 tokens, respectively. Specifically, T2LegalSFG has the highest average input and output length. For each task, we describe the task definition, significance, data acquisition and preprocessing, representative examples, evaluation rubric, and the construction of checklist-mapped references in Appendix B.
3.2 Benchmark Construction Procedure
@@
ExpertLongBench draws data from two primary sources: 6 newly curated task data and 5 adapted existing task data as shown in Tab. 1. The first source includes tasks newly developed by the authors specifically for this benchmark, designed to address gaps in existing resources and better capture real-world expert challenges. Tasks are selected by discussing with domain researchers and practitioners, who also contributed to rubric design. In parallel, we also look into existing datasets and corpora that fit the goal of accessing the capabilities of LLMs on expert-level tasks. To adapt these for our benchmark, we carefully select representative and challenging samples from the original datasets with details and examples shown in Appendix B. Moreover, we design a checklist-based rubric for the tasks and create checklist-mapped references for the samples.
@@
For each task, we design a checklist-based fine-grained evaluation rubric to assess model performance. Our rubric design follows two complementary approaches: 1) Expert-guided design: we collaborate closely with domain experts to co-design evaluation criteria that reflect professional standards and practical requirements and needs. 2) Protocol-refinement design: for tasks with established evaluation protocols guided by experts, we design fine-grained criteria by refining and expanding upon existing standards, ensuring more granular and systematic assessment of model performance. An example rubric showcasing a subset of its checklist items is illustrated in Fig. 1. In appendix B, we show all checklist items for each task. Notably, these rubrics are written by experts, ensuring high-quality criteria that existing LLMs cannot yet replicate (refer to Appendix LABEL:ap:add_discussion). In future work, we will explore methods for automatically generating high-quality checklists. Additional examples along with detailed rubric designs for each task are in the Evaluation Rubric section for each task in Appendix B.
Sample Selection
@@
We establish criteria for further refining the sample pool in tasks with an initial sample size exceeding 100, ultimately selecting 100 representative samples for each task. Our selection criteria focus on two key aspects: diversity and difficulty. For diversity, the select samples encompass a broad range of variations. For example, we select samples covering a wide range of sectors and companies for T10FinanceESG and prioritize the diversity of diagnosis types for T9MedicalDR. For difficulty, we identify key factors influencing difficulty for each task. For instance, we collaborate with legal researchers to determine factors such as document length, number of appeals, number of complaints, and number of dockets for T1LegalMDS. We applied these criteria across most tasks. Details are provided in the Appendix B.
Checklist-mapped Reference Creation
@@
To better evaluate models’ capability for long-form generation on expert-level tasks according to the designed fine-grained rubric, we construct checklist-mapped references by extracting the content corresponding to each checklist item from the human reference using GPT-4o. Exceptions include T3MaterialSEG and T10FinanceESG, where the checklist-based references are constructed during the data collection process (T3) or the human references are well-structured and can simply process to structured checklist-mapped reference (T10). Example checklist-mapped references for T1LegalMDS are shown in Fig. 1 and the checklist-mapped reference subsections in Appendix B.
@@
Specifically, we adopt a role-playing strategy by prompting GPT-4o to extract all relevant information for each checklist item as comprehensively as possible from the reference. If no such information is present, the model is instructed to return “N/A”. Example extraction prompts are illustrated in the checklist-mapped reference subsections in Appendix B. Throughout this paper, we will use the following notation: the checklist for a given task is denoted as $\{c_{i}\}_{i=1}^{n}$, where $n$ is the number of checklist items associated with the specific task. To operationalize this extraction, we design instructions for each checklist item and use GPT-4o to extract the checklist information.
@@
Among the tasks where mapping references to checklists is straightforward, we selected two tasks for both human evaluation and LLM-as-judge evaluation to ensure quality. Evaluation results show that the mapped references achieve over 90% faithfulness and coverage. Detailed information is provided in §4.2. For other tasks, human inspection was employed to ensure the mapping quality. Additional details can be found in Appendix B.

### v3

@@ Abstract
This paper introduces ExpertLongBench, an expert-level benchmark containing 11 tasks from 9 domains that reflect realistic expert workflows and applications. Beyond question answering, the application-driven tasks in ExpertLongBench demand long-form outputs that can exceed 5,000 tokens and strict adherence to domain-specific requirements. Notably, each task in ExpertLongBench includes a rubric, designed or validated by domain experts, to specify task requirements and guide output evaluation. Furthermore, we propose Clear, an evaluation framework that supports accurate evaluation of long-form model outputs in our benchmark. To achieve fine-grained, expert-aligned evaluation, Clear derives checklists from both model outputs and references by extracting information corresponding to items in the task-specific rubric. Checklist items of model outputs are then compared with corresponding items of reference outputs to assess their correctness, enabling grounded evaluation. We benchmark 13 popular large language models (LLMs) and analyze components in Clear, showing that (1) existing LLMs, with the top performer Gemini-2.5-Pro achieving only a 33.4 F1 score, require significant improvement for expert-level tasks; (2) models can generate content corresponding to the required aspects, but far from correct; and (3) accurate checklist extraction and comparison in Clear can be achieved by open-weight models for more scalable, reproducible, and low-cost usage.
@@ 1 Introduction
Large language models (LLMs) have been integrated into applications that demand domain-specific expertise, such as student tutoring (Sonkar et al., 2024), legal case summarization (Siino et al., 2025), and medical diagnosis (McDuff et al., 2025). These expert-level tasks pose greater challenges because they require specialized knowledge and strict adherence to domain standards, where human typically acquires these through advanced education and professional training. Furthermore, many real-world expert-level tasks require understanding lengthy inputs (Wang and Brorsson, 2025) and generating complex and nuanced long-form outputs (Shen et al., 2022; Cascella et al., 2023).
@@
Existing expert-level benchmarks such as MMLU (Hendrycks et al., 2020) and GPQA (Rein et al., 2024) prioritize ease of evaluation of multiple-choice or short-form answers at the expense of alignment with realistic expert applications. Though ExpertQA (Malaviya et al., 2023) assesses longer-form, domain-specific outputs, it remains centered on QA scenarios rather than end-to-end expert workflows. In addition to task scope, a key limitation is the lack of evaluation methods tailored to the specific requirements of each task. WildBench (Lin et al., 2024) creates checklists to evaluate model responses to user queries, yet their model-generated checklists fall short in capturing domain-specific requirements. Finally, a broader challenge is the lack of references in expert-level benchmarks with long-form responses, leading to ungrounded evaluations and missing recall estimation.
@@
To address these gaps, this paper first introduces ExpertLongBench, a multi-disciplinary, expert-level benchmark with tasks that require long-form outputs. ExpertLongBench comprises 11 tasks across 9 domains, with 1050 samples in total (Tab.˜1). Our benchmark enables more comprehensive assessment of a model’s problem-solving capabilities by including domain-specific end-to-end applications, such as drafting legal briefs or clinical notes, with maximum input and output lengths exceeding 200K and 5K tokens, significantly longer than existing datasets (see comparison in Appendix H). The completion of these tasks is time-consuming for experts. For example, it takes over 10 hours for a proficient legal practitioner to summarize a complex legal case, as they need to follow the docket and review tens to hundreds of court filings (Shen et al., 2022). Moreover, ExpertLongBench provides expert-written references per task to support grounded evaluation.
@@
Figure 1: Pipeline of Clear. The example shown is from task T1: multi-document legal case summarization. The checklist mapper takes as input the model output (or human-written reference) and extracts checklist items according to the rubric. Checklists of the model output and the reference are compared at the item level, and the results are subsequently aggregated into the final scores.
@@
To ensure more accurate evaluation of model performance on expert-level tasks, we further propose ClearCheckList-based Expert-level Assessment with Rubric (Fig.˜1). We collaborate with domain experts to develop and validate a detailed evaluation rubric for each task. Unlike subjective criteria, such as “usefulness” (Malaviya et al., 2023) and “helpful information” (Lin et al., 2024), our rubric specifies the essential elements required in the output for each application, enabling fine-grained evaluation that closely align with domain-specific requirements. For example, a well-written summary of legal case documents intended for lawyers must accurately identify the cause of action, the relevant statutory or constitutional basis, and the remedy sought. Based on the rubric, Clear breaks down model responses and references into itemized checklists. For each item, our framework compares the corresponding information extracted from the model’s output against that of the reference, thereby ensuring a grounded and objective checklist-based assessment.  [^1] The benchmark leaderboard and links to code and public data, along with the necessary evaluation resources: https://huggingface.co/spaces/launch/ExpertLongBench.
@@
We evaluate the performance of 13 frontier LLMs on ExpertLongBench, including both open-weight and proprietary model families. Our findings are: $\bullet$ The best model achieves an average F1 score of 33.4 across all tasks, revealing that the end-to-end expert-level tasks in ExpertLongBench present significant challenges for existing LLMs. $\bullet$ Despite a low overall output quality, models can generate content that matches over 67% of the aspects required in the checklist, suggesting a risk of models producing content that appears expert-aligned but is incorrect and thus potentially misleading. $\bullet$ Clear can be configured with open-weight models for cost-effective and scalable benchmarking, as evidenced by a Pearson correlation of 0.88 between the scores assigned by Qwen2.5-72B and GPT-4o for checklist-based evaluation. To ensure ExpertLongBench remains relevant and reliable as LLM capabilities evolve, we commit to continuous development with an interdisciplinary community, incorporating new tasks, high-quality data, and domain-specific evaluation insights.
@@ 2 Related Work
Previous studies have assessed domain-specific knowledge using multiple-choice questions (e.g., MMLU (Hendrycks et al., 2020), GPQA (Rein et al., 2024)) or short-answer questions (e.g., AGIEval (Zhong et al., 2023), SciEval (Sun et al., 2024), HLE (Phan et al., 2025), OlympiadBench (He et al., 2024)). Their scope remains constrained to exam-style questions, failing to reflect the long-form nature of realistic domain-specific tasks. ExpertQA curates open-ended questions written by domain experts (Malaviya et al., 2023). However, the answers average only about 100 words and target information seeking expert knowledge rather than the full scope of an end-to-end workflow. Though DOLOMITES (Malaviya et al., 2025) and ResearchQA (Yifei et al., 2025) target domain-specific long-form generation tasks, they only cover methodological writing and research question answering problems.
ExpertLongBench fills these gaps by introducing diverse, expert-level, long-form tasks that align with complete workflows and features lengthy inputs that represent practical uses; and providing rubric-based evaluation for fine-grained, domain-specific assessment. Further discussions of differences from existing benchmarks are in Appendix H.
Table 1: Benchmark statistics. : rubric is developed by experts; otherwise, it is created by refining and expanding upon established evaluation protocols; : task data is held privately. For each task, we report whether the task data is newly created (✓) or adapted from previous work; the average number of checklist items in each sample (#Rubric); and the average length of the input (#Input) and human reference (#Reference). Several of these tasks feature significantly longer inputs and references compared to existing domain-specific datasets (see Appendix H).
@@
For granular evaluation of a generated text against a reference, recent research explores fact decomposition-based methods, extracting atomic facts from both the texts for comparison using NLI models or LLMs (Min et al., 2023; Kamoi et al., 2023). Nevertheless, the fact decomposition process lacks task-specific considerations for fact granularity, which can result in inconsistent evaluations (Hu et al., 2025).
Another line of work on “LLM-as-a-judge” prompts LLMs to assess generated outputs against general, high-level criteria such as coherence and relevance (Liu et al., 2023). To enhance task specificity, checklist-based evaluations are incorporated into WildBench (Lin et al., 2024) and BiGGenBench (Kim et al., 2024), which evaluate outputs against instance-specific checklists derived from task requirements. Beyond these benchmarks, recent checklist-based methods, e.g., TICK (Cook et al., 2024), CheckEval (Lee et al., 2024a), RocketEval (Wei et al., 2025), and LLM-Rubric (Hashemi et al., 2024), adopt structured rubrics to guide LLM judges. However, these approaches are either not tailored to domain-specific requirements or are not grounded in references, leading to subjective assessments and hindering the accurate evaluation of content coverage and relevance.
@@
3 Benchmark Construction with Multi-disciplinary Expert Tasks
Overview.
We introduce ExpertLongBench, a multi-disciplinary benchmark designed to evaluate the capabilities of LLMs on real-world expert-level tasks that require long-form inputs and outputs. ExpertLongBench features a fine-grained evaluation rubric tailored to reflect the rigorous standards and nuanced requirements of expert domains. The benchmark covers 11 expert-level tasks as shown in Tab.˜1, including six tasks with newly collected data featuring unique challenges such as super-long input. Notably, rubric design is complex and highly time-intensive; for instance, creating the rubric for T1LegalMDS required over 10 hours of expert effort. Details regarding the experts involved in rubric design and verification are provided in Appendix A. The selection of these tasks adhere to specific criteria: (1) it is possible to create well-defined rubric for consistent and objective evaluation; (2) it requires domain expertise for both task resolution and assessment; and (3) it is grounded in real-world expert workflows. Additional details are discussed in Appendices A and B. Generally, each sample in ExpertLongBench includes the following key elements:
  • Task Input: The context of the task. For example, the inputs for T2LegalSFG consist of an average of 14 long transcripts used to generate a Statement of Fact.
  • Human-written Reference: Each task includes human-authored references in natural language. Details can be referred to Appendix B.
  • Checklist-mapped Reference: A checklist-based, fine-grained evaluation rubric is designed for each task by collaboration with domain experts and referencing established resources. A checklist-mapped reference is constructed for every sample, as detailed in §3.
Tab.˜1 summarizes key statistics of our benchmark. In total, it contains 1,050 samples, with an average input length and human reference length of 36,204 and 851 tokens, respectively. For each task, we describe the task definition, significance, data acquisition and preprocessing, representative examples, evaluation rubric, and the construction of checklist-mapped references in Appendix B.
While public evaluation sets enable transparency and facilitate model development, they are also prone to contamination and overfitting. We design our benchmark with both a public and private subset as shown in Tab.˜1. The public set supports open experimentation, while the private set contains sensitive data used for fair evaluation to ensure robustness and confidentiality. [^2] We provide a submission channel on the benchmark website that allows researchers to evaluate models on the private subset. The public set is shared under the CC BY-NC-SA 4.0 license [^3] https://creativecommons.org/licenses/by-nc-sa/4.0/, while the private set remains confidential. Details are provided in Appendix I.
@@ 3 Benchmark Construction with Multi-disciplinary Expert Tasks
ExpertLongBench draws data from two primary sources: 6 newly curated task data and 5 adapted existing task data. The first source includes tasks newly developed by the authors to address gaps in existing resources and better capture real-world expert challenges. Tasks are selected by discussing with domain researchers and practitioners, who also contributed to rubric design. In parallel, we also look into existing datasets and corpora that fit the goal of accessing the capabilities of LLMs on expert-level tasks. To adapt these for our benchmark, we carefully select representative and challenging samples with details and examples shown in Appendix B. Moreover, we design a checklist-based rubric for the tasks and create checklist-mapped references.
@@
For each task, we design a checklist-based fine-grained evaluation rubric that is applicable to all data points within the same task to assess model performance. Our rubric design follows two complementary approaches: 1) Expert-guided design: we collaborate closely with domain experts [^4] Experts involved in designing or verifying the rubrics possess relevant academic or professional expertise, with details provided in Appendix A. to co-design evaluation criteria that reflect professional standards and practical requirements and needs. 2) Protocol-refinement design: for tasks with established evaluation protocols guided by experts, we design fine-grained criteria by refining and expanding upon existing standards, ensuring more granular and systematic assessment of model performance. An example rubric showcasing a subset of its checklist items is illustrated in Fig.˜1. In Appendix B, we show all checklist items for each task. Notably, these rubrics are written by experts, ensuring high-quality criteria that existing LLMs cannot yet replicate (refer to Appendix I), pointing to future research on automatically generating high-quality checklists. Additional examples along with detailed rubric designs for each task are in the Evaluation Rubric section for each task in Appendix B.
Sample Selection.
@@
We establish criteria for further refining the sample pool in tasks with an initial sample size exceeding 100, ultimately selecting 100 representative samples for each task. Our selection criteria focus on two key aspects: diversity and difficulty. For diversity, the select samples encompass a broad range of variations. For difficulty, we identify key factors influencing difficulty for each task. For instance, we determine factors such as document length, number of appeals, number of complaints, and number of dockets for T1LegalMDS. We applied these criteria across most tasks. Details are provided in the Appendix B.
Checklist-mapped Reference Creation.
@@
To evaluate models’ capability for long-form generation on expert-level tasks according to the designed fine-grained rubric, we construct checklist-mapped references by extracting the content corresponding to each checklist item from the reference using GPT-4o. Exceptions include T3MaterialSEG and T10FinanceESG, where the checklist-based references are constructed during the data collection process (T3) or the references are already well-structured and can be processed into a checklist format (T10). Example checklist-mapped references for T1LegalMDS are shown in Fig.˜1 and the checklist-mapped reference subsections in Appendix B.
@@
Specifically, we adopt a role-playing strategy by prompting GPT-4o to extract all relevant information for each checklist item as comprehensively as possible from the reference. If no such information is present, the model is instructed to return “N/A” (prompts are in checklist-mapped reference subsections in Appendix B). Throughout this paper, we will use the following notation: the checklist for a given task is denoted as $\{c_{i}\}_{i=1}^{n}$, where $n$ is the number of checklist items associated with the specific task. We design instructions for each checklist item and use GPT-4o to extract the corresponding information.
@@
We then use both human evaluation and LLM-as-a-judge evaluation to ensure the quality of the extraction. Evaluation results show that the mapped references achieve over 90% faithfulness and coverage on two tasks. Detailed information is provided in §4.2. For other tasks, human inspection was employed to ensure the mapping quality. Additional details can be found in Appendix B.